OpenStudy (anonymous):

Find an equation in standard form of the parabola passing through the given points. (0,3).(1,2),(2,3)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

how far did you get?

OpenStudy (anonymous):

I don't know where to even start.

jimthompson5910 (jim_thompson5910):

what is the general equation for a parabola?

jimthompson5910 (jim_thompson5910):

or a quadratic

OpenStudy (anonymous):

Like the quadratic equation? \[x=\frac{-b \pm \sqrt{b^2-4ac} }{ 2a }\]

jimthompson5910 (jim_thompson5910):

kinda, but I was aiming more towards y = ax^2 + bx + c

jimthompson5910 (jim_thompson5910):

look familiar?

OpenStudy (anonymous):

Yes sorry.

jimthompson5910 (jim_thompson5910):

so we have one point (0,3) which means x = 0, y = 3 plug them into y = ax^2 + bx + c y = ax^2 + bx + c 3 = a(0)^2 + b(0) + c 3 = a*0 + b*0 + c 3 = 0 + 0 + c 3 = c c = 3

jimthompson5910 (jim_thompson5910):

What do you get when you plug (1,2) into y = ax^2 + bx + c

OpenStudy (anonymous):

2=a+b+c

jimthompson5910 (jim_thompson5910):

good, since c = 3, we can plug that in as well 2=a+b+c 2=a+b+3 2-3=a+b -1 = a+b a+b = -1

jimthompson5910 (jim_thompson5910):

From there, we can isolate one variable. Let's pick on b a+b = -1 b = -1-a b = -a - 1

jimthompson5910 (jim_thompson5910):

What do you get when you plug (2,3) into y = ax^2 + bx + c

OpenStudy (anonymous):

3=4a+2b+c a=1

OpenStudy (anonymous):

So y=x^2-2x+3

jimthompson5910 (jim_thompson5910):

let me check

jimthompson5910 (jim_thompson5910):

3=4a+2b+c 3=4a+2b+3 ... plug in c = 3 3=4a+2(-a - 1)+3 ... plug in b = -a - 1 3=4a-2a - 2+3 3=2a + 1 3-1 = 2a 2 = 2a 2a = 2 a = 2/2 a = 1 so you are correct

jimthompson5910 (jim_thompson5910):

b = -a - 1 b = -1-1 b = -2 also correct

jimthompson5910 (jim_thompson5910):

so yes, y = x^2 - 2x + 3

jimthompson5910 (jim_thompson5910):

I recommend checking that each point lies on this parabola.

OpenStudy (anonymous):

Ok thanks.

jimthompson5910 (jim_thompson5910):

you're welcome